

********************************************************************
UARS WEEKLY STATUS REPORT - 04 NOV 2005
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2005099 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004274 and 2005099 are waiting for removal of these files by the DAAC,
indicating completion of the archiving process.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2005099. MEPS data between 2004260 and
2005099 are waiting for removal of these files by the DAAC, indicating
completion of the archiving process.
HEPSB data files are version 1. HEPSB data files between 1991274 and
1991366 have been processed and are archived at the GSFC DAAC. HEPSB data
files between 1991274 and 1991336 are waiting for removal of these files by
the DAAC, indicating completion of the archiving process.
Processing of all PEM DAAC data files has been suspended until the
DAAC catches up with archiving all processed data files. An error was
discovered in the DAAC IP which caused problems with DAAC access of PEM
Level 2 data. This was corrected at SwRI and the DAAC was successfully
able to access SwRI. SwRI was waiting for the DAAC to catch up on Level 2
PEM files. However, during this process, the DAAC again ran into access
problems. It was found that the DAAC IP node had changed again which
disrupted DAAC access of PEM Level 2 data. Investigation concluded that
the DAAC at GSFC changed network locations and PEM is now in the approval
process to update the DAAC IP access address.
Checks on the PEM data availability through the DAAC have been conducted
by SwRI. These tests were prompted by a user of PEM data who has had trouble
accessing the PEM data in the DAAC. The documentation link to the PEM HEPSA
Level 2 Document still points to the old version file pem_l2_hepsa_v01.txt
instead of the v02 file for hepsa data. Also, there is no MEPS entry for the
MEPS data set. These problems have been reported to help@daac.gsfc.nasa.gov
in mid-October.
Tests of the AXIS1 data file formats have been successful. The AXIS1
format design has been coded. A test file has been generated. A check
program with processing scripts have also been generated. Example binary
read program is complete and verified. These results have been sent to
the project. Creation of the AXIS1 data documents for DAAC data will
begin shortly.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered since updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG is participating in the on-going UARS end of mission
planning activities. Operation of UARS instruments is
scheduled to end with the turn-off of HALOE on November 21,
2005.
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 22 October 2005 (Orbit 77260, GMT Day 295) through
28 October 2005 (Orbit 77367, GMT Day 301).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
-NONE THIS REPORT PERIOD
A Reverse-to-Forward Yaw Around maneuver (number 144) was successfully
executed in Orbit 77301 beginning 297/ 16:42:55Z and completing at
297/ 17:20:05Z on 24 October 2005. All instruments were in their off
or standby mode. Due to the SA position at 265 degrees, the effective
night length of the spacecraft after the maneuver was approximately
52.43 minutes. Post maneuver analysis of data indicate that during
the Yaw Around, the end-of-night load bus voltage was 26.24 volts and
the average depths of discharge (DODs) on Batteries 2 and 3 were
approximately 11.88 and 23.42 percent respectively. The next (forward-
to-reverse) YAW is scheduled for 23 November 2005 (DOY 327).
The beta angle ranged between the angles of 10.1 degrees to 13.3
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 40 degrees on 09 November 2005 (DOY 313).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. TMON 45 and 46 are
maintaining the relay operations for battery 2. Battery 2 and Battery
3 are currently maintaining greater than 23.6 V end-of-night (EON)
load bus voltage. Battery temperatures are stable with a temperature
delta between Battery 2 and Battery 3 of 3.12 to 2.87 degrees C.
Additional battery performance data for each battery on 21 October
2005 and 28 October 2005 are provided in Appendix D.
The clock error ranged from 10.7 to -10.6 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 04
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 02
Four (4) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continues to closely monitor the condition of
battery 2 and battery 3. Battery 2 and 3 are sharing load a bit more
now that the adjustment to TMON45 was made. The FOT is very pleased
with the current operational mode of the batteries and feel confident
in our ability to support upcoming HALOE instrument operations. NASA
power branch has been providing close support in monitoring the UARS
battery condition. The FOT has been authorized to disable all
platform and instrument heaters on the SSPP. This will be done at the
earliest available opportunity if needed.
While performing post-burn analysis of the 20 October 2005 burn, the
FOT noted that the FOT calculated amount of fuel remaining was almost
20 KG less than the amount that FDF had anticipated. Based on the FOT
calculated amount, there was not enough fuel remaining onboard to
execute another 18-minute burn. FDF had estimated enough fuel left
for 2 more complete 18-minute burns with fuel remaining to do a short
3-5 minute burn. The FOT and FDF are working together to find an
explanation for the difference. The FOT reviewed pre-launch and early
launch documentation to provide initial conditions for FDF and to
ensure that the FOT method used for determining fuel remaining was
correct. The FOT is also attempting to verify the values for the
constants used in the calculation.
MEETINGS
The FOT attended a debrief meeting on 25 October 2005. Results of the
burn on 20 October 2005 were presented by FDF and the FOT.
The FOT also attended a Debris Avoidance Working Group (DAWG) meeting
on 26 October 2005. The topic discussed was the UARS / ISS conjunction
analysis being done by FDF and Cheyenne Mountain. The FOT proposed
doing our last burns out of phase with ISS to allow enough time for
proper conjunction analysis to be done by ISS and Cheyenne Mountain
once the last UARS burn is done in December.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 23 November 2005 (DOY 327).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
The FOT has been given direction by NASA Program Manager, Julio
Marius, to complete all decommissioning activities and passivate UARS
by 15 December 2005.
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 28 October 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
77279 296/0541 05-378 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
77292 297/0145 05-379 MPS VT L5 TO VT L6
77301 297/1655 05-380 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
77301 297/1655 05-379 MACS YAW MANEUVER
77361 301/1325 05-381 OBC/CLK CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 28 October 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 28 October 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
77279 296/0545 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
77301 297/1659 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
77361 301/1329 CLK ADJ FR 0.0 TO -20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 28 October 2005
21 October 2005 - GMT Day 294 Beta = 15.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -2.42/ -3.68 +00.0/+00.0 04.6 0.00 0.0
2 -700.0/-700.0 +3.23/ +2.60 +23.2/-12.0 26.1 1.06 11.1
3 224.0/ 00.0 +5.79/ +4.83 +31.2/-22.8 25.8 1.02 22.9
28 October 2005 - GMT Day 301 Beta = 13.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -2.11/ -3.05 +00.0/+00.0 04.6 0.00 0.0
2 -700.0/-700.0 +3.87/ +2.92 +21.6/-09.8 26.4 1.07 9.3
3 005.6/ -72.8 +7.09/ +5.79 +32.0/-21.2 26.1 1.02 23.6
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
77292 297/0145 LV5 -> LV6
APPENDIX E
UARS ANOMALY REPORTS
Week ending 28 October 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-053 FOT 77307 298/0100 41213 DATA Dropouts
05-054 HGA 77337 300/0006 41233 HGA Continuity Error
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-052 FOT 77237 293/1241 41080 Post-burn late Acq-2:43
recov
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-042 MPS 76339 233/2347 N/A BATTERY 2 DIFFV -700mV
05-047 FOT See List Oct 05 30832 Generic Late Acquisition
05-053 FOT 77307 298/0100 41213 DATA Dropouts
05-054 HGA 77337 300/0006 41233 HGA Continuity Error